The normal body-temperature of a person is 97°F. Calculate the rate at which heat is flowing out of his body through the clothes assuming the following values. Room temperature = 47°F, surface of the body under clothes = 1.6 m2, conductivity of the cloth = 0.04 J/m-s-°C, thickness of the cloth = 0.5 cm.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. K = 0.04 J/m-5°C, A = 1.6 m^2 T base 1 = 97°F = 36.1°C t base 2 = 47°F = 8.33°C l = 0.5 cm = 0.005 m Q/t = KA(θ base 1 – θ base 2)/l = 4 * 10^-2 * 1.6 * 27.78/5 * 10^-3 = 356 J/s